# balnet: Pathwise Estimation of Covariate Balancing Propensity Scores

### v1

@@ Abstract
We present balnet, an R package for scalable pathwise estimation of covariate balancing propensity scores via logistic covariate balancing loss functions. Regularization paths are computed with Yang and Hastie (2024)’s generic elastic net solver, supporting convex losses with non-smooth penalties, as well as group penalties and feature-specific penalty factors. For lasso penalization, balnet computes a regularized balance path from the largest observed covariate imbalance to a user-specified fraction of this maximum. We illustrate the method with an application to spatial pixel-level balancing for constructing synthetic control weights for the average treatment effect on the treated, using satellite data on wildfires.
@@ 1 Introduction
A primary use of the propensity score is inverse probability weighting (IPW), where outcomes are reweighted to construct treated and control groups that are comparable in terms of their pre-treatment covariates. Under the true (oracle) propensity score, IPW exactly balances covariate means between treatment arms and the overall population. However, even without model-misspecification, this exact balancing property does not generally hold for propensity scores estimated using standard maximum likelihood logistic regression. Maximum likelihood estimation prioritizes accurate prediction of treatment assignment, rather than covariate balance. Imai and Ratkovic (2014) introduced the notion of covariate balancing propensity scores, propensity score estimators explicitly designed to optimize covariate balance. Causal effects estimated using these kinds of weights are more robust to model mis-specification, and more efficient than approaches that estimate propensity scores via maximum likelihood (Ben-Michael et al., 2021; Chattopadhyay et al., 2020).
@@
Our package, balnet for R (R Core Team, 2024), addresses these limitations by combining covariate balancing loss functions with modern optimization techniques for large-scale regularization problems. Our approach uses proximal quasi-Newton methods for generic convex losses, including algorithmic refinements for coordinate descent such as screening and active-set rules (Friedman et al., 2010; Simon et al., 2013; Tibshirani et al., 2012). The core algorithm builds on the adelie C++ elastic net solver by Yang and Hastie (2024), interfaced with R via Rcpp (Eddelbuettel and François, 2011), and relies on numerical linear algebra routines from Eigen (Guennebaud et al., 2010). In an application from Wu et al. (2023) involving spatial balancing with about 140,000 observations and 500 covariates, balnet computes the full regularization path, from raw, unweighted imbalance down to commonly accepted balance thresholds, in under one minute on a standard laptop.
@@ 2 Regularization paths for covariate balancing propensity scores
where $\eta_{i}=\beta_{0}+X_{i}\beta$. This can be verified by setting the gradient of $l_{1}(\eta)$ to zero, which yields the balance conditions in (4). The IPW estimator constructed using such weights has favorable statistical properties: when the logistic model is correctly specified it is $\sqrt{n}$-consistent with efficient variance (Wager, 2024, Theorem 7.1). Intuitively, this approach targets the ratio $W_{i}/e(X_{i})$ (the Riesz representer for $\mathbb{E}\left[Y_{i}(1)\right]$; see e.g. Hirshberg and Wager (2021)), rather than estimating $e(X_{i})$ via maximum likelihood and inverting the fitted probabilities [^2] These weights can also improve doubly robust methods by stabilizing the AIPW debiasing term (Robins et al., 1994); see, for example, Vermeulen and Vansteelandt (2015)..
@@
Remark 1.
For estimating the ATE, $\mathbb{E}\left[Y_{i}(1)\right]-\mathbb{E}\left[Y_{i}(0)\right]$, this approach requires fitting two propensity score models: one targeting balance for the treated mean and one for the control mean, unlike maximum likelihood which fits one set of coefficients.
Remark 2.
For software, it is sufficient to implement only the loss in (5). By symmetry, propensities targeting $\mathbb{E}\left[Y_{i}(0)\right]$ are obtained by solving (5) with inverted treatment indicators $W_{i}^{\prime}=1-W_{i}$. The propensities targeting $\mathbb{E}\left[Y_{i}(0)\right]$ can also be used to target the ATT, $\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid W_{i}=1\right]$, since the losses only differ by a scaling (the former weights the control to match $\frac{1}{n}\sum_{i=1}^{n}X_{i}$ while the latter weights the controls to match $\frac{1}{\sum_{i=1}^{n}W_{i}}\sum_{i=1}^{n}W_{i}X_{i}$).
In practice, directly minimizing (5) may not be feasible. Exact balance may be unattainable in settings with high-dimensional covariates ($p>n$) or limited overlap, potentially producing extreme IPW weights. The covariate balancing loss imposes stronger requirements than maximum likelihood estimation, which only requires overlap, e.g., the absence of complete separation.
@@ 3 Application: Pixel-level balancing
We focus on a single focal year (2008) and land cover type (conifer forests), yielding a dataset with $n=141,780$ observations (4,466 treated) and $p=553$ covariates, measuring past fire history as well as topographic, meteorological, disturbance, and vegetation characteristics. The goal is to find weights that match the means of all 553 covariates in the synthetic control group to the corresponding treated-group means. The largest standardized mean difference in the unweighted data is approximately 470, occurring for snow water equivalent in September 2004. We set the parameter max.imbalance to 0.05, targeting a maximum allowable standardized covariate imbalance of 0.05. Internally, balnet adjusts the $\lambda$ sequence to range from $\lambda^{\max}=470$ to $\lambda_{\min}=0.05$ over nlambda =100=100 logarithmically spaced values.
@@
Let $\bar{X}_{\text{target}}$ and $\bar{\sigma}_{\text{target}}$ denote the treated-group covariate means and standard deviations, and let
@@ 3.1 Timings
To assess runtime, we conduct a small timing experiment using the data from this section [^6] We do not include benchmarks against the R packages mentioned in Section 1, as under our experimental setup they either returned errors or did not complete within a one-hour time window.. We define a baseline dataset consisting of the first n=100,000n=100{,}000 observations and $p=500$ covariates from the data described in the previous section. We then double both the sample size and the number of covariates by sampling with replacement from the baseline data. For each setting, we measure the time required for balnet to compute a regularization path targeting the ATT with maximum imbalance $\lambda_{\min}\in\{0.05,0.01\}$. Table 1 reports runtimes obtained on a 8-core laptop with 24 GB of memory, using four cores for solver parallelism. The results suggest that runtime is more sensitive to the $\lambda$-sequence endpoints than to increases in the number of observations and covariates.
@@
Table 1: Runtimes (in seconds) and relative scaling when doubling both the number of observations and covariates, and tightening balance requirement.
@@ 4 Discussion
Regularization is a practical necessity in many modern statistical settings and pathwise solutions are now a standard tool in regularized statistical learning (Friedman et al., 2007, 2010). We use pathwise regularization in a causal setting to target balancing weights for treatment effects under unconfoundedness using logistic covariate balancing loss functions. A natural direction for future work is panel-data settings where separate balancing weights are estimated along both time and unit dimensions, with regularization paths defined accordingly.

### v2

@@ Abstract
We present balnet, an R package for scalable pathwise estimation of covariate balancing propensity scores via logistic covariate balancing loss functions. Regularization paths are computed with Yang and Hastie (2024)’s generic elastic net solver, supporting convex losses with non-smooth penalties, as well as group penalties and feature-specific penalty factors. For lasso penalization, balnet computes a regularization path of balancing weights from the largest observed covariate imbalance to a user-specified fraction of this maximum. We illustrate the method with an application to spatial pixel-level balancing for constructing synthetic control weights for the average treatment effect on the treated, using satellite data on wildfires.
@@ 1 Introduction
A primary use of the propensity score is inverse probability weighting (IPW), where outcomes are reweighted to construct treated and control groups that are comparable in terms of their pre-treatment covariates. Under the true (oracle) propensity score, IPW exactly balances covariate means between treatment arms and the overall population. However, even without model misspecification, this exact balancing property does not generally hold for propensity scores estimated using standard maximum likelihood logistic regression, which prioritizes accurate prediction of treatment assignment, rather than covariate balance. Imai and Ratkovic (2014) introduced the notion of covariate balancing propensity scores, propensity score estimators explicitly designed to optimize covariate balance. Causal effects estimated using these balancing weights can be more robust to model misspecification, and more efficient than approaches that estimate propensity scores via maximum likelihood (Ben-Michael et al., 2021; Chattopadhyay et al., 2020). These weights can also improve doubly robust methods, such as AIPW (Robins et al., 1994), by stabilizing the debiasing term (van der Laan et al., 2024; Vermeulen and Vansteelandt, 2015).
@@
Our package, balnet for R (R Core Team, 2024), addresses these limitations by combining covariate balancing loss functions with modern optimization techniques for large-scale regularization problems. Our approach uses proximal quasi-Newton methods for generic convex losses, including algorithmic refinements for coordinate descent such as screening and active-set rules (Friedman et al., 2010; Simon et al., 2013; Tibshirani et al., 2012). The core algorithm builds on the adelie C++ elastic net solver by Yang and Hastie (2024), interfaced with R via Rcpp (Eddelbuettel and François, 2011), and relies on numerical linear algebra routines from Eigen (Guennebaud et al., 2010). In an application from Wu et al. (2023) involving spatial balancing with about 140,000 observations and 500 covariates, balnet computes the full regularization path, from raw, unweighted imbalance down to commonly accepted balance thresholds, within minutes on a standard laptop.
@@ 2 Regularization paths for covariate balancing propensity scores
where $\eta_{i}=\beta_{0}+X_{i}\beta$. This follows by setting the gradient of $l_{1}(\eta)$ to zero, which yields the balance conditions in (4). The IPW estimator constructed using such weights has favorable statistical properties: when the logistic model is correctly specified it is $\sqrt{n}$-consistent with efficient variance (Wager, 2024, Theorem 7.1). An equivalent perspective is that we target the Riesz representer for $\mathbb{E}\left[Y_{i}(1)\right]$, the ratio $W_{i}/e(X_{i})$, rather than estimating $e(X_{i})$ via maximum likelihood and inverting the fitted probabilities (e.g., Hirshberg and Wager (2021); see also Chernozhukov et al. (2022)).
@@
For estimating the ATE, $\mathbb{E}\left[Y_{i}(1)\right]-\mathbb{E}\left[Y_{i}(0)\right]$, balnet fits two propensity score models: one targeting balance for the treated mean and one for the control mean, unlike maximum likelihood which fits one set of coefficients [^2] For software, it is sufficient to implement only the loss in (5). By symmetry, propensities targeting $\mathbb{E}\left[Y_{i}(0)\right]$ are obtained by solving (5) with inverted treatment indicators $W_{i}^{\prime}=1-W_{i}$. The propensities targeting $\mathbb{E}\left[Y_{i}(0)\right]$ can also be used to target the ATT, $\mathbb{E}\left[Y_{i}(1)-Y_{i}(0)\mid W_{i}=1\right]$, since the losses only differ by a scaling (the former weights the control to match $\frac{1}{n}\sum_{i=1}^{n}X_{i}$ while the latter weights the controls to match $\frac{1}{\sum_{i=1}^{n}W_{i}}\sum_{i=1}^{n}W_{i}X_{i}$)..
In practice, directly minimizing (5) may not be feasible. Exact balance may be unattainable in settings with high-dimensional covariates ($p>n$) or limited overlap, potentially producing extreme weights. The covariate balancing loss imposes stronger requirements than maximum likelihood estimation, which only requires overlap, e.g., the absence of complete separation.
@@ 3 Application: Pixel-level balancing
We focus on a single focal year (2008) and land cover type (conifer forests), yielding a dataset with $n=141,780$ observations (4,466 treated) and $p=553$ covariates, measuring past fire history as well as topographic, meteorological, disturbance, and vegetation characteristics. The goal is to find weights that match the means of all 553 covariates in the synthetic control group to the corresponding treated-group means. The largest standardized mean difference in the unweighted data is approximately 470, occurring for snow water equivalent in September 2004. We set the parameter max.imbalance to 0.05, targeting a maximum allowable standardized covariate imbalance of 0.05. Internally, balnet adjusts the $\lambda$ sequence such that the resulting maximum standardized imbalance ranges from 470 down to the target 0.05 over nlambda =100=100 logarithmically spaced values.
fit <- balnet(X, W, target = "ATT", max.imbalance = 0.05,
                verbose = TRUE, num.threads = 4)
@@
balnet provides extractor methods to seamlessly retrieve these weights for any $\lambda$ on the path. Let $\bar{X}_{\text{target}}$ and $\bar{\sigma}_{\text{target}}$ denote the treated-group covariate means and standard deviations, and let
@@ 3.1 Timings
To assess runtime, we conduct a small timing experiment using the data from this section [^6] We do not include benchmarks against the R packages mentioned in Section 1 as they did not produce valid results under our experimental setup: CBPS did not complete within a one-hour time window, RCAL  returned results for only the first value of the regularization grid, with the remaining 99 entries being NA, and the remaining packages crashed on initialization.. We define a baseline dataset consisting of the first n=100,000n=100{,}000 observations and $p=500$ covariates from the data described in the previous section. We then double both the sample size and the number of covariates by sampling with replacement from the baseline data. For each setting, we measure the time required for balnet to compute a regularization path targeting the ATT with maximum imbalance $\lambda_{\min}\in\{0.05,0.01\}$. Table 1 reports runtimes obtained on an 8-core laptop with 24 GB of memory, using four cores for solver parallelism. The results suggest that runtime is more sensitive to the $\lambda$-sequence endpoints than to increases in the number of observations and covariates.
@@
Table 1: Runtimes (in seconds) and relative scaling when doubling both the number of observations and covariates, and tightening balance requirement. The number of cores used is 4.
@@ 4 Discussion
Regularization is a practical necessity in many modern statistical settings and pathwise solutions are now a standard tool in regularized statistical learning (Friedman et al., 2007, 2010). We use pathwise regularization in a causal setting to target balancing weights for treatment effects under unconfoundedness using logistic covariate balancing loss functions. A natural direction for future work is to extend this approach to additional causal estimands and to panel-data settings where separate balancing weights are estimated along both time and unit dimensions, with regularization paths defined accordingly.
A direction for future algorithmic work is improved tolerance scaling near the end of the regularization path, where full balance may be unattainable and the solver must truncate. Deviance-based scaling, standard in canonical GLM solvers, is not directly applicable here: developing analogous criteria tailored to balancing losses could reduce runtimes. A simple heuristic is to relax the coordinate descent tolerance parameter tol, trading solution accuracy for faster path computation.
Acknowledgments
We are grateful to Brandon de la Cuesta and Rina Park for helpful feedback.